Find the smallest positive root for the given function by using the bisection method with accuracy 10^-3 f(x) = 2x5 – x3
Answer #1

find the smallest positive root for the given benetion by bisection method with accuracy of 103. the using f(x) = 225-23 Forf(X.)L0; f(x3770 Iteration 3 .. 24 -1-0.5 2 =X+23 = 2 E- 0.75 .: f(24) = f(-0.75) = -0.0527 LO f(X₂) >0 and f (24) Lo IteratiIteration 9 f(x8) 70 and f(g) 20 i. x 10 0.703125 -0.7109375 = 88+Xg 2 2 0.70703125 f(x,0) = f(-0.70703125) = 7.5562970 Itera